A roller coaster is traveling upside down in a vertical loop of diameter 25m. What is its minimum speed if the riders are not about to fall out of their seats at the highest point in the loop?

*Please be careful with significant figures.
Answer #1

For the roller coaster not to fall , the centripetal force at the highest point is equal to the weight or the body.

So( mv​​​​​2)/r = mg thus v = √gr

v = √ 25×9.8 = 15.65 m/s.